OpenStudy (anonymous):

Newton's method to find all roots of the equation correct to six decimal places. (Enter your answers as a comma-separated list.) 6 cos x = x + 1

OpenStudy (tkhunny):

And "Newton's Method" would be... What? \(f(x) = 6\cos(x) - x - 1\) The challenge is to find \(f(x) = 0\) \(f'(x) = -6\sin(x) - 1\) Okay, now build it!

OpenStudy (anonymous):

\[x-((6\cos(x)-x-1) \div (-6\sin(x)-1))\] isnt that it

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

oh your still here cool

OpenStudy (tkhunny):

Be sure. Is it \(\dfrac{f(x)}{f'(x)}\;or\;\dfrac{f'(x)}{f(x)}\)?

OpenStudy (anonymous):

f(x)/f'(x)

OpenStudy (anonymous):

i got that part, but then i dont really understand the process that comes after

OpenStudy (anonymous):

why should i start off with an initial value of 1, why not 0?

OpenStudy (tkhunny):

Okay, now we need to know where to start. What do you suggest? The amplitude of the cosine is 6 and the slope of the line is +1, so I would suspect that ALL roots are in [-7,5], otherwise the linear function is outside the Range of the Cosine. Make sense?

OpenStudy (anonymous):

well not really how did you get -7,5

OpenStudy (anonymous):

i understand the amp is 6 and slope is 1

OpenStudy (tkhunny):

It's a linear function, x+1. For what values does that land in [-6,6], where the cosine lives?

OpenStudy (anonymous):

oh ok well wouldnt that be -5,7

OpenStudy (tkhunny):

No. -5 + 1 = -4 7 + 1 = 8 Yes -7 + 1 = -6 5 + 1 = 6

OpenStudy (anonymous):

ohhhhh ok

OpenStudy (anonymous):

so do i plug those into my formula

OpenStudy (tkhunny):

Now, to your question. You do NOT want to start anywhere the derivative is very close to zero. This will shoot you off somewhere and possible never come back. Look what an initial x = 0 does. 0 5 5.904173 6.994311 6.292835 5.070488 It immediately goes past out limit of 5 and stays out there until it goes insane and starts jumping all over. f'(0) = -1 -- That could be steeper. f'(1) = -6.049 -- Oh, I like that!

OpenStudy (anonymous):

so i have to look at the derivative formula and imagine what number would be far enough away from zero?

OpenStudy (anonymous):

because if the derivative is zero its a flat line?

OpenStudy (tkhunny):

Fair enough. A quick sketch of f'(x) wouldn't hurt. A derivative of zero is very bad news. That will never work out.

OpenStudy (anonymous):

but one is around half way between -7 and 6?

OpenStudy (tkhunny):

So? Stay away from it. It looks like -6.5, -3, 0, 3, and 6 would be bad places to start. Start somewhere else.

OpenStudy (tkhunny):

Start at x = 1. What do you get?

OpenStudy (anonymous):

1.24181

OpenStudy (anonymous):

wait no sorry

OpenStudy (tkhunny):

Should be 1.205298

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

so now i assighn that to x in my calculator ??

OpenStudy (tkhunny):

Yes. And the next iteration...

OpenStudy (tkhunny):

Note: I'm not rounding intermediate results to 6 decimal places. My results may differ a little.

OpenStudy (anonymous):

1.19608

OpenStudy (tkhunny):

Okay. I got 1.196090, but that's fine. Probably one or two more iterations.

OpenStudy (anonymous):

that means keep pushing enter on my calculator? lol

OpenStudy (tkhunny):

That is one of the advantages of Newton's Method. It is "Self-Correcting". If we're off by some tiny amount of rounding, the method will just fix the error. It's pretty sweet. Maybe. It depends on how your function is set up. If you can do that, let's see the next two iterations.

OpenStudy (anonymous):

i was supposed to get 3 answers i guess, i guess the only way i could know there were 3 roots would be to graph it?

OpenStudy (tkhunny):

No, that is not the only way. I like to refer to "Blanket Bombing". We created a Domain. Why not use all of it? Hunt around a little?

OpenStudy (tkhunny):

-7 leads to a failure to converge -6 leads to the middle root -5 leads to the left-most root -4 leads to the left-most root -3 leads to a failure to converge -2 leads to the middle root -1 leads to the middle root 0 leads to a failure to converge 1 leads to the right-most root 2 leads to the right-most root 3 leads to the middle root 4 leads to the left-most root 5 leads to a failure to converge You need to be satisfied that you found all of them. I find behavior around 2, 3, and 4 a little peculiar. I might be feeling a need to investigate around there a little more before I was completely satisfied. Really, though, with this nice little investigation, we keep tripping over the same three roots. Time to call it solved and move on.